# Machine Unlearning: A Comprehensive Survey

### v1

@@ 7. Privacy Threats on Machine Unlearning
Although machine unlearning was first proposed to protect users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously. Similar to machine learning scenarios, there are two main kinds of privacy threats in unlearning. They are membership inference attacks and private information reconstruction attacks. In this section, we introduce the two prevalent unlearning privacy threats detailedly.
7.1. Membership Inference Attacks in Unlearning
@@
Chen et al. (Chen et al., 2021b) first pointed out that when a model is unlearned, the difference in the outputs from the model before and after unlearning leaks the privacy of the erased data. Then they proposed the corresponding membership inference attack pipeline in unlearning, which includes three phases: posteriors generation, feature construction, and membership inference.
@@
    Posteriors Generation. Suppose that the attacker has the ability to access two versions of the ML model, the model $\theta_{\mathcal{A}}$ before unlearning and the unlearned model $\theta_{\mathcal{U}}$. Assume a target sample $e$, the attacker queries $\theta_{\mathcal{A}}$ and $\theta_{\mathcal{U}}$, and obtains the corresponding posteriors, $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, which also called as confidence values in (Shokri et al., 2017).
@@
    Feature Construction. After achieving the two posteriors $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, the attacker sums them to make the inference feature vector $F$. Common methods exist to construct the feature vector shown in (Chen et al., 2021b).
@@
    Inference. After the attacker finishes training the attack model based on the created features $F$, he inputs the collected feature to the inference model to predict if the specific sample $e$ is in the erased dataset of unlearning models.
@@
In (Chen et al., 2021b), they assumed that the attacker has admission to two versions of ML models before and after unlearning, but it is sometimes impractical, especially in black-box learning scenarios. Lu et al. (Lu et al., 2022) further proposed a label-only membership inference attack method to infer if a data point is unlearned by the model, eliminating the dependence on accessing posteriors. Their basic idea is that the same noise injection on candidate data points will show different results for the sample in or not in the training dataset. Thus, they made the adversary continuously query the original and unlearned models and add noise to change their outputing labels. By observing the disturbance amplitude, they can determine whether an item is deleted.
@@
Golatkar et al. (Golatkar et al., 2020b) derived an upper bound to confirm the maximizing knowledge that can be extracted from a black-box model. They queried the model with an image and obtained the related output. They then used the entropy of the result probabilities to construct an effective black-box membership inference (Shokri and Shmatikov, 2015) attack in machine unlearning.
7.2. Privacy Reconstruction Attacks in Unlearning
@@
Privacy reconstruction is another popular attacking kind in machine unlearning. In an unlearning scenario, Gao et al. (Gao et al., 2022) proposed the deleted reconstruction attacks to recover the removed data from the outputs of the original and unlearning models. In their work, they formalized erasure inference and erasure reconstruction attacks, in which the attacker aims to either infer which record is deleted or reconstruct the erased sample. In particular, for the deletion inference, they formalize the goal of an erasure inference to distinguish a data example $e$ was in or not in the erased dataset, $e\in D_{e}$ or $e^{\prime}\notin D_{e}$. For the deletion reconstruction, they focused on reconstructing the erased example $e$. In all their reconstruction attacks, the attacker is not given any particular examples, and its goal is to extract knowledge about the features of the erased instance. Specifically, the deletion reconstructions include the deleted instance reconstruction and deleted label reconstruction. As named, the deleted instance reconstruction is to extract all of the features of the erased instance, and the erased label reconstruction is to infer the label of the erased point in the classification problem.
@@
Zanella-Béguelin et al. (Zanella-Béguelin et al., 2020) indicated that the releasing snapshot of overlapped language models would leak the privacy of the training dataset. They verified that the model updates pose a significant risk to the private information added to or removed from the training dataset by a huge amount of experimental results. Zanella-Béguelin et al. found five phenomenons. First, an attacker can extract particular sentences used or not in the training dataset by comparing two models. Second, analysing more model snapshots shows more information about the updated data than considering fewer model snapshots. Third, adding or deleting other non-private data during model updates can not mitigate privacy leakage. Fourth, differential privacy can reduce privacy leakage risks, but it also decreases the accuracy of trained models. Fifth, to mitigate the privacy leakage risks while keeping the model utility, the server can limit the model parameters access or only output a subset of the results.
@@
Many studies furthermore utilised these privacy threats to evaluate the unlearning effect. Huang et al. (Huang et al., 2021a) proposed Ensembled Membership Auditing (EMA) for auditing data erasure. They use the membership inference to evaluate the removing effect of unlearning. Graves et al. (Graves et al., 2021) indicated that if an attacker can infer the sensitive information that was wanted to be erased, then it means that the server has not guarded the user’s rights to be forgotten. Baumhauer et al. (Baumhauer et al., 2022) developed linear filtration to sanitise classification models with logits prediction after class-wide deletion requests. They verified their methods by testing how well it defends against privacy attacks.
@@
Both (Chen et al., 2021b) and (Zanella-Béguelin et al., 2020) pointed out that differential privacy guarantees that a model does not reveal too much information about any particular training data point. A differentially private model can furthermore guarantee the group’s privacy by bounding the impacts of a group of training samples. If using DP to protect the privacy of a group of $|D\backslash D_{e}|$training samples against snapshot attacks on $\theta_{D},\theta_{D\backslash D_{e}}$, it means that $\theta_{D\backslash D_{e}}$ cannot be more useful than $\theta_{D}$.
@@ 8. Machine Unlearning Applications
Literature Application Scenarios Algorithms Realization Methods Evaluation Metric Year
(Huang et al., 2021b) Data unlearnable EMP Add noise ACC 2021
@@
(Du et al., 2019) Anomaly detection Unlearning Maintain memory set ##\##FP and ##\##FN 2019
@@
  • ##\##FP: number of false positive, ##\##FN: number of false negative
@@
Besides the inherent demands in machine unlearning that draw much research attention, many researchers also find that machine unlearning can be applied in many other scenarios to solve related problems. We list the recent unlearning applications in Table 3. The most popular application of machine unlearning is to mitigate the anomaly, including the backdoor triggers and pollution, from a trained model.
@@
Wang et al. (Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs. They designed two approaches to eliminate backdoor-related neurons/weights from the infected model and repair the infected model to be robust against malicious images. Then, Liu et al. (Liu et al., 2022b) proposed that a backdoor model learned the poisoned decision boundary. Data points with triggers are all classified into the target class. They reversed the backdoor injection process to defend against it in machine unlearning, which is simple but effective. Their method contains two primary steps: first, they use a max-entropy staircase approximator to complete trigger reconstruction; second, they remove the added backdoor triggers using unlearning. They named these two key steps of BAERASER as trigger pattern recovery and trigger pattern unlearning. Via a dynamic penalty mechanism, they mitigated the sharp accuracy degradation of gradient-ascent-based machine unlearning methods.
@@
Repairing pollution is another successful unlearning application. Cao et al. (Cao et al., 2018) proposed KARMA to search different subsets of training datasets and return the subset that with the highest misclassifications. KARMA includes three parts. First, KARMA searches for potential causality that leads to the wrong classification of a ML model. It clusters the misclassified data points into different parts and extracts the centres of clusters. Then KARMA prioritizes the search for similar data samples in the training datasets based on these extracted centres. Second, KARMA grows the causality found in the first stage by finding more training samples and forming a cluster. Third, KARMA determines if a causality cluster is polluted and calculates how many samples the cluster contains.
@@
Machine unlearning was also applied to lifelong anomaly detection problems. Du et al. (Du et al., 2019) proposed new approaches to solve the challenges of exploding loss and sharp accuracy degradation caused by unlearning. In their presentation, they explain their techniques, emphasizing handling false negatives. When meeting such case $x_{t}$, their goal is, therefore, to modify the model to reduce the probability of $p(x_{t})$, or to increase the loss $\mathcal{L}(x_{t})$.
@@
Huang et al. (Huang et al., 2021b) presented a method that can make training examples unlearnable by injecting a type of error-minimizing noise. This kind of noise is intentionally generated to reduce the error of the training examples close to zero, which can mislead the model into considering there is ”nothing” to learn from these examples. They first tried the error-maximizing (Goodfellow et al., 2014) noise but found that this noise cannot prevent DNN learning when applied in a sample-wise manner to the training data points. Therefore, they then begin to study the opposite direction of error-maximizing noise. In particular, they proposed the error-minimizing noise to stop the model from being punished by the loss function during traditional ML model training. Therefore, it can mislead the ML model to consider that there is ”nothing” to learn.
@@ 9. Lessons Learnt and Discussions
In the research domain of machine unlearning, researchers mainly face three difficult challenges, the stochasticity of training, incrementality of training, and catastrophe of unlearning, when they try to design efficient and effective unlearning mechanisms replacing naive retraining. They tried many mechanisms to mitigate the influence of these challenges. For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models. In approximate unlearning, they bounded the removed estimation to aviod the accuracy degradation. However, it is still not easy to analyze them clearly and solve them totally. Moreover, as the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met new challenges. In this section, we will discuss the differences between various unlearning scenarios and their corresponding challenges. Based on these challenges, we list some potential research directions on which we can focus.
In the common centralized unlearning scenario, retraining from scratch can achieve the best unlearning effect, but it is expensive in both computation and storage. Existing methods try to design new unlearning mechanisms to reduce the cost from the two aspects. Although they proposed many methods, they just mitigated the influence of the main challenges and still have not solved them. Existing unlearning methods were at the beginning step that tried to implement effective unlearning. They proposed mechanisms, but most of them cannot guarantee the final results, so they also bound the unlearned item (Guo et al., 2020) or set an unlearning threshold (Nguyen et al., 2020). Although they also introduced many metrics to verify unlearning effect, the unlearning and verification processes are split. It means that the unlearning result is uncertain during the unlearning period before they finish the verification. Therefore, we find and list some open questions in centralized unlearning.
@@
    Optimizing the original three challenges. The first two challenges are hard to solve, so recent researches focus on reducing the utility degradation after unlearning. Even though many studies exist, there is still a long way to eliminate the unlearning degradation.
@@
    How to unlearn with a certain or exact goal? In other words, can we unlearn as learning? We know the purpose, to what extent we have unlearned and when we can finish or stop unlearning process.
Machine unlearning in distributed scenarios has many differences to which in centralized scenarios. We take federated unlearning as a representative example of distributed unlearning. The first difference is that federated unlearning can only be implemented locally on the client’s side if they want to unlearn some specific samples because clients’ do not upload their data to the FL server in a federated scenario. To avoid interacting with clients during unlearning, researchers (Liu et al., 2021; Wu et al., 2022b) proposed to unlearn the contribution of a whole client while not some samples of the client. The second difference is that when unlearning requests come during the FL training process, the FL server must first execute the unlearning process and broadcast the unlearned model for later updating to avoid other clients wasting computation on the before-unlearned model. The third difference is that federated unlearning is more vulnerable to catastrophic degradation than centralized unlearning because if the FL server broadcasts the catastrophic unlearned model and other clients update based on the unlearned model, it will vanish the efforts of other clients that trained before. After introducing these differences, we can see that the challenges in federated unlearning are more complex than in centralized unlearning, and we conclude the following open problems in federated unlearning.
@@
    Federated unlearning cannot use the fast retraining methods because data is out of reach for the server. Therefore, federated unlearning can only be implemented using approximate unlearning methods. However, as we know, approximate unlearning is easy to cause catastrophic unlearning, and federated learning is more vulnerable to degradation, so how to control the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
@@
    Privacy in federated learning is more sensitive than in other situations because FL was designed to protect participants’ privacy at the beginning. May unlearning bring other privacy demands in federated learning, such as secret or unnoticeable unlearning?
@@
Besides exploring machine unlearning based on regularly structured data, researchers tried to implement unlearning in graph data. Exact unlearning may be suitable for centralized graph unlearning if graph data is sparse. However, the challenges of approximate unlearning may be more difficult than structured-data-based unlearning because, in graph unlearning, the relationship and influence between data samples are more complex than structured data (Chen et al., 2021a; Chien et al., 2022). In particular, graph data includes not only the node feature value but also the connecting edge information. Therefore, in graph unlearning, the estimation of the contribution of a node will be more difficult than in regular data unlearning. The original problems in regular data unlearning will be more challenging in graph unlearning. Besides these problems, graph unlearning also face unique problems that is related to edge structure information. In graph unlearning, how to unlearn some edges or a sub-graph is a big question.
@@
Another important part is the privacy and security issues in machine unlearning. Although machine unlearning was proposed to protect users’ privacy when they want to leave, it brings new threats in that adversaries have a chance to infer the information about the removed data. Literature (Chen et al., 2021b; Zanella-Béguelin et al., 2020) have pointed out that updates of unlearning will leak privacy information, and they proposed corresponding attacks to infer this private information. However, most recent unlearning privacy leakage attacks are similar to those in a learning situation. An attack that is tailored to unlearning mechanisms is expected. A similar situation exists in unlearning applications. Although researchers proposed to unlearn a backdoor trigger (Liu et al., 2022b) or pollution (Cao et al., 2018), they only used a few unlearning techniques and paid more effort to detect those anomalies. One important reason is that unlearning mechanism is not mature enough. We are at the beginning of machine unlearning, and there are still many problems in unlearning itself that need to be solved. Finding machine unlearning applying situation and tailoring unlearning techniques to this situation is the direction of unlearning application.

### v2

@@ 7. Privacy and Security Threats on Machine Unlearning
Although unlearning methods were initially proposed to safeguard users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously. Similar to machine learning scenarios, there are two main kinds of privacy threats in unlearning. They are membership inference attacks and private information reconstruction attacks. In this section, we detail the two prevalent unlearning privacy threats.
7.1. Privacy Threats on Machine Unlearning
7.1.1. Membership Inference Attacks in Unlearning
@@
Chen et al. (Chen et al., 2021) first pointed out that when a model is unlearned, the discrepancy in the outputs from the model before and after unlearning leaks privacy of erased data. Then, they proposed the corresponding membership inference attack pipeline in unlearning, which includes three phases: posterior generation, feature construction, and membership inference.
@@
    Posteriors Generation. Suppose that the attacker can access two versions of the trained model, the model $\theta_{\mathcal{A}}$ before unlearning and the model $\theta_{\mathcal{U}}$ after unlearning. Assume a target data point $e$, the attacker queries $\theta_{\mathcal{A}}$ and $\theta_{\mathcal{U}}$, and has the corresponding posteriors, $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, which also called as confidence values in (Shokri et al., 2017).
@@
    Feature Construction. After achieving the two posteriors $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, the attacker sums them to make the inference feature vector $F$. Common methods exist to construct the feature vector shown in (Chen et al., 2021).
@@
    Inference. After the attacker finishes training the attack model based on the created features $F$, he inputs the collected feature to the inference model to predict if the specific sample $e$ occurs in the erased dataset of unlearning models.
@@
In (Chen et al., 2021), they assumed that the attacker has admission to two ML models before and after unlearning, but it is sometimes impractical, especially in black-box learning scenarios. Lu et al. (Lu et al., 2022) further proposed a label-only membership inference method to imply if a sample is unlearned, eliminating the dependence on accessing posteriors. Their basic idea is that the same noise injection on candidate data points will show different results for the sample in or not in the training dataset. Thus, they made the adversary continuously query the original and unlearned models and add noise to modify their outputs. Observing the disturbance amplitude lets them determine whether an item is deleted.
@@
Golatkar et al. (Golatkar et al., 2020b) derived an upper bound to confirm the maximizing knowledge that can be extracted from a black-box model. They queried the model with a picture and obtained the related output. They used the entropy of the result probabilities to construct an effective black-box membership inference (Shokri and Shmatikov, 2015) attack in machine unlearning.
7.1.2. Privacy Reconstruction Attacks in Unlearning
@@
Privacy reconstruction is another popular attack in machine unlearning. In an unlearning scenario, Gao et al. (Gao et al., 2022) proposed the deleted reconstruction attacks to recover the removed data from the outputs of the original and unlearning models. In their work, they formalized erasure inference and erasure recovering attacks. The attacker seeks to infer which sample is removed or recovers the erased sample. In particular, for the deletion inference, they formulate the objective of an erasure inference to decide if a data instance $e$ was in or not in the erased dataset, $e\in D_{e}$ or $e^{\prime}\notin D_{e}$. For the deletion reconstruction, they focused on reconstructing the erased example $e$. In all their reconstruction attacks, the attacker does not have any particular samples, and the purpose is to extract the features knowledge of the erased example. Specifically, the deletion reconstructions include the deleted instance reconstruction and deleted label reconstruction. As named, the deleted instance reconstruction is to extract all of the information of the erased example, and the erased label reconstruction is to infer the label of the erased point in the classification problem.
@@
Zanella-Béguelin et al. (Zanella-Béguelin et al., 2020) indicated that the releasing snapshot of overlapped language models would leak the privacy of the training dataset. They verified that the model updates significantly threaten the private information added to or deleted from the training dataset by many experimental results. Zanella-Béguelin et al. found five phenomenons. First, an attacker can extract particular sentences used or not in the training dataset by comparing two models. Second, analyzing more model snapshots shows more information about the updated data than considering fewer model snapshots. Third, adding or deleting other non-private data during model updates can not mitigate privacy leakage. Fourth, differential privacy can reduce privacy leakage risks and decrease trained models’ accuracy. Fifth, to mitigate the privacy leakage risks while keeping the model utility, the server can limit the model parameters access or only output a subset of the results.
@@
Many studies further utilized these privacy threats to evaluate the unlearning effect. Huang et al. (Huang et al., 2021a) proposed Ensembled Membership Auditing (EMA) for auditing data erasure. They use the membership inference to assess the removing effectiveness of unlearning. Graves et al. (Graves et al., 2021) indicated that if an attacker can infer the sensitive information that was wanted to be erased, then it means that the server has not guarded the rights to be forgotten. Baumhauer et al. (Baumhauer et al., 2022) developed linear filtration to sanitize classification models with logits prediction after class-wide deletion requests. They verified their methods by testing how well the method defends against privacy attacks.
@@
Both (Chen et al., 2021) and (Zanella-Béguelin et al., 2020) pointed out that differential privacy guarantees that a model does not reveal too much knowledge about any training sample. A DP-protected model can further guarantee the group’s privacy by binding the impacts of a bunch of training samples. If using DP to protect the privacy of a bunch of $|D\backslash D_{e}|$training examples against snapshot attacks on $\theta_{D},\theta_{D\backslash D_{e}}$, it means that $\theta_{D\backslash D_{e}}$ cannot be more useful than $\theta_{D}$.
7.2. Security Threats on Machine Unlearning
There are several papers that address security threats targeting machine unlearning. For instance, Marchant et al. (Marchant et al., 2022) examine an attacker’s strategy to increase the computational cost of performing machine unlearning. Studies by Zhao et al. (Zhao et al., 2024) and Hu et al. (Hu et al., ) investigate security vulnerabilities within unlearning systems, specifically by uploading customized malicious data update requests to negatively influence the model utility. Additionally, some studies propose backdooring or poisoning methods that exploit unlearning requests to achieve backdoor insertion, as discussed by (Di et al., 2022) and (Liu et al., 2024). These works collectively highlight the need for robust security measures in machine unlearning to mitigate these emerging threats.
@@ 8. Machine Unlearning Applications
Literature Application Scenarios Methods Realization Methods Evaluation Metric Year
(Huang et al., 2021b) Data “unlearnable” EMP Add noise ACC 2021
@@
(Du et al., 2019) Anomaly detection - Maintain memory set ##\##FP and ##\##FN 2019
@@
  • ##\##FN: number of false negative, ##\##FP: number of false positive
@@
Besides the inherent demands of machine unlearning that draw much research attention, many researchers also find that machine unlearning can be applied in many other scenarios to solve related problems. We list the recent unlearning applications in Table 5. The most popular application of machine unlearning is to mitigate the anomalies, including backdoor triggers and pollution from an already-trained model.
@@
Study (Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs. They designed two approaches to eliminate backdoored neurons from the backdoored model and repair the model to be strong against malicious pictures. Then, in (Liu et al., 2022b), they offered that a backdoor model learned the poisoned decision boundary. Data points with triggers are all classified into the target class. They reversed the backdoor injection process to defend against it in machine unlearning, which is simple but effective. Their method contains two primary steps: first, they utilize a max-entropy staircase approximator to complete trigger reconstruction; second, they remove the added triggers using unlearning. They named these two key steps of BAERASER as trigger pattern recovery and trigger pattern unlearning. Via a dynamic penalty mechanism, they mitigated the sharp accuracy degradation of gradient-ascent-based machine unlearning methods.
@@
Repairing pollution is another successful unlearning application. Cao et al. (Cao et al., 2018) proposed “KARMA” to search various subsets of original datasets and return the subset with the highest misclassifications. First, KARMA searches for possible reasons that lead to the wrong ML model classification. It clusters the misclassified samples into various domains and extracts the middle of clusters. KARMA prioritizes the search for matching examples in the original datasets using these extracted centers. Second, KARMA grows the reason discovered in the first step by discovering more training samples and creating a cluster. Third, KARMA determines if a causality cluster is polluted and calculates how many samples the cluster contains.
@@
Machine unlearning was also applied to lifelong anomaly detection problems. Du et al. (Du et al., 2019) proposed new approaches to solve the challenges of exploding loss and sharp accuracy degradation caused by unlearning. In their presentation, they explain their techniques, emphasizing handling false negatives. When meeting such case $x_{t}$, their purpose is, therefore, to modify the model to reduce the possibility of $p(x_{t})$, or to increase the loss $\mathcal{L}(x_{t})$.
@@
Huang et al. (Huang et al., 2021b) presented a method that can make samples unlearnable by injecting error-minimizing noise. This noise is intentionally synthesized to diminish the error of the samples close to zero, which can mislead the model into considering there is ”nothing” to learn. They first tried the error-maximizing noise but found that it could not prevent DNN learning when used sample-wise to the training data points. Therefore, they then begin to study the opposite direction of error-maximizing noise. In particular, they proposed the error-minimizing noise to stop the model from being punished by the loss function during traditional ML model training. Therefore, it can mislead the ML model to consider that there is ”nothing” to learn.
@@ 9. Lessons Learnt and Discussions
In the research domain of machine unlearning, researchers mainly face three difficult challenges, the stochasticity of training, incrementality of training, and catastrophe of unlearning, when they try to design efficient and effective unlearning mechanisms replacing naive retraining. They tried many mechanisms to mitigate the influence of these challenges. For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models. In approximate unlearning, they bounded the removed estimation to avoid accuracy degradation. However, it is still not easy to analyze them clearly and solve them totally. Moreover, as the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met corresponding new challenges. Here, we discuss the differences between various unlearning scenarios and their corresponding challenges. Based on these challenges, we list some potential research directions on which we can focus.
In the common centralized unlearning scenario, retraining from scratch can achieve the best unlearning effect, but it is expensive in both computation and storage. Existing methods try to design new unlearning mechanisms to reduce the cost from the two aspects. Although they proposed many methods, they just mitigated the influence of the main challenges and still have not solved them. Existing unlearning methods were at the beginning phase, which is trying to implement effective unlearning. They proposed mechanisms, but most of them cannot guarantee the final results, so they also bound the unlearned item (Guo et al., 2020) or set an unlearning threshold (Nguyen et al., 2020). Although they also introduced many metrics to verify the unlearning effect, the unlearning and verification processes are split. This means that the unlearning result is uncertain during the unlearning period before the verification is finished. Therefore, we find and list some open questions in centralized unlearning.
@@
    Optimizing the existing challenges. The first two challenges are hard to solve exactly, so recent research focuses on reducing utility degradation after unlearning. Even though many studies exist, there is still a long way to go to eliminate the degradation of unlearning.
@@
    How do we unlearn with a certain or exact goal? In other words, can we unlearn as learning? If we know the purpose, to what extent we have unlearned, and when we can finish or stop the unlearning process.
Machine unlearning in distributed scenarios has many differences from centralized scenarios. We take federated unlearning as a representative example of distributed unlearning. The first difference is that federated unlearning can only be implemented locally on the client’s side if they want to unlearn some specific samples because clients do not upload their data to the FL server in a federated scenario. To avoid interacting with clients during unlearning, researchers (Liu et al., 2021; Wu et al., 2022b) proposed to unlearn the contribution of a whole client while not some samples of the client. The second difference is that when unlearning requests come during the FL training process, the FL server must first execute the unlearning process and broadcast the unlearned model for later updating to avoid other clients wasting computation on the before-unlearned model. The third difference is that federated unlearning is more vulnerable to catastrophic degradation than centralized unlearning because if the FL server broadcasts the catastrophic unlearned model and other clients update based on the unlearned model, it will vanish the efforts of other clients that trained before. After introducing these differences, we can see that the challenges in federated unlearning are more complex than in centralized unlearning, and we conclude the following open problems in federated unlearning.
@@
    Federated unlearning cannot use the fast retraining methods because data is out of reach for the server. Therefore, federated unlearning can only be implemented using approximate unlearning methods. However, as we know, approximate unlearning easily causes catastrophic unlearning, and federated learning is more vulnerable to degradation, so controlling the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
@@
    Existing federated unlearning methods require the activation of all users, including those without unlearning requests, to assist in the unlearning process. These approaches are impractical and inefficient, particularly when unlearning requests are frequent. Therefore, it is crucial to study how to balance the unlearning effect and efficiency in federated unlearning.
@@
Besides exploring machine unlearning based on regularly structured data, researchers tried to implement unlearning in graph data. Exact unlearning may be suitable for centralized graph unlearning if graph data is sparse. However, the challenges of approximate unlearning may be more difficult than structured-data-based unlearning because, in graph unlearning, the relationship and influence between data samples are more complex than structured data (Chen et al., 2022b; Chien et al., 2022a). In particular, graph data includes not only the node feature value but also the connecting edge information. Therefore, in graph unlearning, the estimation of the contribution of a node will be more difficult than in regular data unlearning. The original problems in regular data unlearning will be more challenging in graph unlearning. Besides these problems, graph unlearning also faces unique problems that are related to edge structure information. In graph unlearning, unlearning some edges or sub-graphs is a big question.
After designing machine unlearning algorithms, effective verification and auditing methods are necessary (Thudi et al., 2022b). Most existing unlearning verification methods rely on backdooring techniques. However, these methods inherently degrade model utility because they require mixing backdoored samples into the model training process. Investigating ways to preserve model utility in backdoor-based unlearning verification methods is an under-explored area. Additionally, backdoor-based verification methods must include backdoored samples in the unlearning requests to verify the effectiveness of these requests. This requirement restricts these methods from supporting verification for single-sample unlearning requests. Exploring new strategies that can effectively verify unlearning requests without compromising model utility or being suitable for both single-sample and multi-sample unlearning scenarios remains a significant challenge in this field.
@@
Another important part is the privacy and security issues in machine unlearning. Machine unlearning was first proposed to protect users’ privacy, but it brings new threats in that adversaries have a chance to infer the information about the removed data. Literature (Chen et al., 2021; Zanella-Béguelin et al., 2020) have pointed out that updates of unlearning will leak privacy information, and they proposed corresponding attacks to infer this private information. However, the most recent unlearning privacy leakage attacks have been similar to those in a learning situation. An attack that is tailored to unlearning mechanisms is expected. A similar situation exists in unlearning applications. Although researchers proposed to unlearn a backdoor trigger (Liu et al., 2022b) or pollution (Cao et al., 2018), they only used a few unlearning techniques and paid more effort to detect those anomalies. One important reason is that the unlearning mechanism is not mature enough. We are at the beginning of machine unlearning investigation, and there are still many under-explored problems in unlearning itself. Finding machine unlearning applying situation and tailoring unlearning techniques to this situation is the direction of unlearning application.